OpenStudy (anonymous):

Let R be the region between the graph of y=e^-2x and the x-axis for x greater than or = to 3. The area of R is?

jimthompson5910 (jim_thompson5910):

are you familiar with integrals?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

its just i dont now what my upper and lower limit should be.

OpenStudy (anonymous):

and im bad at integrating

jimthompson5910 (jim_thompson5910):

Since x is greater than or equal to 3, this means that the lower limit is 3 The upper limit will be infinity. This is valid because e^(-2x) is a decreasing function (it will slowly approach 0)

jimthompson5910 (jim_thompson5910):

So this problem deals with improper integrals. Have you had experience working with improper integrals?

OpenStudy (anonymous):

oh yeah i did

OpenStudy (anonymous):

i dont really remeber it though

jimthompson5910 (jim_thompson5910):

int( e^(-2x) dx) .... Note: this means "integral of e^(-2x) with respect to x" Let u = -2x, so du/dx = -2 ---> du = -2dx ---> dx = -du/2 So int( e^(-2x) dx) turns into int( e^u (-du/2)) Rearrange terms to get -1/2*int( e^u du) and then integrate -1/2*(e^u + C) -1/2*e^u + C -1/2*e^(-2x) + C So the integral of e^(-2x) is -1/2*e^(-2x) + C With me so far?

OpenStudy (anonymous):

ok lemme just read it..give me a minute

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

yupp i get it so far

jimthompson5910 (jim_thompson5910):

that's great...my apologies, I got distracted Now evaluate the integral at its limits In general, if the limits of this integral were from 'a' to 'b', then int(e^(-2x) dx) = (-1/2e^(-2b) + C) - (-1/2e^(-2a) + C) int(e^(-2x) dx) = -1/2e^(-2b) + 1/2e^(-2a) ------------------------------------------------------- In our case, a = 3 and b = t where t approaches infinity, so... int(e^(-2x) dx) = -1/2e^(-2t) + e^(-2*3) int(e^(-2x) dx) = -1/2e^(-2t) + 1/2e^(-6) From here, you take the limit as t goes to infinity. This will make -1/2e^(-2t) approach 0. This effectively makes the first term -1/2e^(-2t) go away leaving us with 1/2e^(-6) This means that the answer is 1/2e^(-6) which can be written as 1/(2e^6) Hopefully this makes sense. If not, let me know. Thanks.

OpenStudy (anonymous):

So u know how u subtract the two values u get when u subsitute the upper and lower limit. When u substitute a 3 u get 1/2e^(-6) and when u subtistute a t it just goes to 0. So the asnwer is just 1/2e^(-6). Am i understanding it correctly?

jimthompson5910 (jim_thompson5910):

when I substitute t and let t approach infinity, the first term -1/2e^(-2t) goes to zero, so yes you have the right idea.

OpenStudy (anonymous):

THANK YOU SO MUCH! AND TAKING THE TIME TO EXPLAIN THOROUGHLY. This was extremely helpful. =).